# Non-singular and probability measure-preserving actions of infinite permutation groups

### v1

@@ 1. Introduction
Roelcke precompact groups share some tameness properties with compact groups. For example, the unitary representations of non-archimedean Roelcke precompact groups can be completely classified [Tsankov2012]. However, in contrast to the situation with compact groups, irreducible representations are usually infinite-dimensional. This classification has proved helpful for studying their probability measure-preserving (pmp, for short) actions. For example, an analogue of de Finetti’s theorem holds for appropriate actions of such groups ([Jahel2022], also see below) and Jahel and Joseph [Jahel2023p] have been able to classify the invariant random subgroups of some of them.
@@
Let $G$ be a non-archimedean, Roelcke precompact, Polish group and let $G\curvearrowright(\Omega,\mu)$ be a non-singular action. Then $\Omega$ can be represented as a disjoint countable union $\bigsqcup\Omega_{i}$ of invariant subsets on each of which the action is isomorphic to an induced action from a pmp action of an open subgroup of $G$. In particular, there exists a $\sigma$-finite measure equivalent to $\mu$ which is preserved by $G$.
@@
Then any $G$-invariant, ergodic measure on $[0,1]^{M}$ is a product measure.
Condition missing (ii) above is satisfied, for example, if the action $G\curvearrowright M^{2}$ has finitely many orbits but also in many other situations. This condition has been considered in the pseudo-finite case in [Liebeck2010]. Both conditions missing (i) and missing (ii) are remnants of compactness and are automatically satisfied if $G$ is Roelcke precompact. Condition missing (i) is also satisfied, for example, if $M$ is a homogeneous structure whose age is given by a first-order universal theory (or, equivalently, a collection of finite forbidden configurations) and $G=\operatorname{Aut}(M)$. An important class of examples where the hypotheses are not satisfied but the conclusion still holds is the rational Urysohn space and some generalizations of it. In a very recent preprint [Barritault2024pp], Barritault, Jahel, and Joseph manage to classify all unitary representations of its isometry group and deduce, in particular, a version of Section 1.
@@ Acknowledgments
I would like to thank Tomás Ibarlucía and Yuri Neretin for bringing to my attention Nessonov’s article [Nessonov2020], which inspired Section 1. I am also grateful to Rémi Barritault, Colin Jahel, and Matthieu Joseph for useful remarks on a preliminary version of the paper.
@@ 2.1. The lattice of algebraically closed sets
missing (i) Write $H=G_{D}$. Let $S_{0},S_{1},\ldots$ be an enumeration of the orbits of $G$ on $M$. Let $A_{i}=\bigcup_{j\leq i}S_{j}\cap D$. As $D$ is locally finite, each $A_{i}$ is finite. Moreover, $G_{A_{0}}\supseteq G_{A_{1}}\supseteq\cdots$ and $\bigcap_{i}G_{A_{i}}=H$. As $H$ is open, by Roelcke precompactness, each $G_{A_{i}}$ contains only finitely many $H$-double cosets, so there is $i$ such that $G_{A_{i}}=H$. Let $A=A_{i}$. We check that $\operatorname{acl}A=D$. The $\subseteq$ inclusion follows from the fact that $D$ is algebraically closed. For the other, if $a\in D$, then $G_{A}\cdot a=H\cdot a\subseteq D\cap G\cdot a$, which is finite by the local finiteness of $D$.
missing (ii) If $gG_{A}g^{-1}\cap G_{A}$ has finite index in $G_{A}$, then $G_{A}g\cdot A$ is finite. Thus $g\cdot A\subseteq\operatorname{acl}A=D$, which implies that $g\cdot A\subseteq A$. As $A$ is finite, this means that $g\cdot A=A$ and thus $g\in G_{A}=G_{D}$.
missing (iii) This follows directly from missing (ii). ∎
@@
missing (i) This follows from [Evans2016]*Lemma 2.4, noting that for $K\leq G$ and $A\in\operatorname{Fin}(M)$, the orbit $K\cdot A$ is infinite iff there exists $a\in A$ with $K\cdot a$ infinite.
missing (ii) Note that $H\lhd G_{D}$ and that $G_{D}/H$ acts faithfully on $D$ with finite orbits. It follows from [Evans2016]*Lemma 2.9 that the action $G_{D}/H\curvearrowright D$ is topologically faithful, so $G_{D}/H$ is a profinite group. As every $G$-orbit splits into finitely many $G_{D}$-orbits by Roelcke precompactness, it suffices to show that every $G_{D}$-orbit splits into finitely many $H$-orbits. Let $A\in\operatorname{Fin}(M)$ and $F\subseteq G_{D}$ be finite such that $G_{D}=HF(G_{A}\cap G_{D})$. Then $F\cdot A$ is a set of representatives for all $H$-orbits in $G_{D}\cdot A$. ∎
@@
At the level of stabilizers, $A\leq B$ iff $G_{A}\cap G_{B}$ has finite index in $G_{B}$, and $A\sim B$ if $G_{A}$ and $G_{B}$ are commensurate. It follows from (missing) 2.1 that for all $A,B\in\operatorname{Fin}(M)$,
@@
Let $G$ be a non-archimedean, Roelcke precompact, Polish group and let $G\curvearrowright M$ be a continuous action on a discrete, countable set. Let the quasi-order $\leq$ be defined as in (missing) 2.1. Then the following hold:
@@
    The relation $<$ on $\operatorname{Fin}(M)$ is well-founded.
@@
missing (i) Let $A,B\in\operatorname{Fin}(M)$. The least upper bound $[A]\vee[B]$ is given by $[A\cup B]$. That $[A\cup B]$ is an upper bound follows from (missing) 2.2. If $[C]\geq[A],[B]$, then $A\cup B\subseteq\operatorname{acl}C$, so $\operatorname{acl}(A\cup B)\subseteq\operatorname{acl}C$ by Subsection 2.1.
For the greatest lower bound, let $D=\operatorname{acl}A\cap\operatorname{acl}B$ and note that $D$ is algebraically closed and locally finite. It is also definable because $G_{(A\cup B)}$ is a subgroup of $G_{D}$. Use Subsection 2.1 to find a finite $D_{0}\subseteq D$ with $D=\operatorname{acl}D_{0}$ and $G_{D}=G_{D_{0}}$. We show that $[D_{0}]=[A]\wedge[B]$. It is clear that $[D_{0}]\leq[A],[B]$. If $C\in\operatorname{Fin}(M)$ is such that $C\leq A,B$, then $C\subseteq\operatorname{acl}A\cap\operatorname{acl}B$, so $\operatorname{acl}C\subseteq\operatorname{acl}A\cap\operatorname{acl}B=\operatorname{acl}D_{0}$, and it follows that $C\leq D_{0}$.
missing (ii) Suppose, towards a contradiction, that there exist $A_{0},A_{1},\ldots\in\operatorname{Fin}(M)$ with $A_{0}>A_{1}>\cdots$. Let $D_{i}=\operatorname{acl}A_{i}$ and note that $D_{i}\supseteq D_{i+1}$ and $A_{i}\subseteq D_{j}$ for all $i\geq j$. Let $H_{i}=G_{(D_{i})}$, so that $H_{0}\leq H_{1}\leq\cdots$.
We prove that for every $i$, the inclusion $H_{i}\cdot A_{0}\subseteq H_{i+1}\cdot A_{0}$ is proper. By assumption, $G_{A_{i+1}}\cdot A_{i}$ is infinite, so $H_{i+1}\cdot A_{i}$ is also infinite by Subsection 2.1 missing (i). On the other hand, $A_{i}\subseteq\operatorname{acl}A_{0}$, so $G_{A_{0}}\cdot A_{i}$ is finite. Hence there exists $g\in H_{i+1}$ with $g^{-1}\cdot A_{i}\notin G_{A_{0}}\cdot A_{i}$. We will show that $g\cdot A_{0}\notin H_{i}\cdot A_{0}$. Suppose, to the contrary, that there is $h\in H_{i}$ with $h\cdot A_{0}=g\cdot A_{0}$. Then $g^{-1}h\in G_{A_{0}}$ and $g^{-1}h\cdot A_{i}=g^{-1}\cdot A_{i}$, contradicting the fact that $g^{-1}\cdot A_{i}\notin G_{A_{0}}\cdot A_{i}$.
We obtained that $H_{0}\cdot A_{0}\subsetneq H_{1}\cdot A_{0}\subsetneq\cdots$ is an infinite strictly increasing sequence of $H_{0}$-invariant subsets of $G\cdot A_{0}$, which contradicts Subsection 2.1 missing (ii). ∎
@@ 2.2. The universal action and unitary representations
Let $G$ be a Roelcke precompact, non-archimedean, Polish group. Among all actions of $G$ on a countable set, there is a universal one that can be constructed as follows. Let $(V_{i}:i\in I)$ be a collection of representatives for the equivalence relation of conjugacy of open subgroups of $G$. Note that the set $I$ is countable by [Tsankov2012]*Corollary 2.5. Let
@@
and equip it with the left translation action. Note that for any open $V\leq G$, there exists $a\in\mathcal{U}$ with $V=G_{a}$ and if $V=N_{G}(V)$, this $a$ is unique. This implies, in particular, that the action $G\curvearrowright\mathcal{U}$ is topologically faithful (see [Evans2016]*Lemma 1.9).
@@
The uniqueness of $a$ is guaranteed by Subsection 2.1. It is easy to check that $A\sim s(A)$ for all $A\in\operatorname{Fin}(\mathcal{U})$ and that $s$ is idempotent in the sense that $s(\{s(A)\})=s(A)$ for all $A\in\operatorname{Fin}(M)$. Thus the set
@@
To see the $\subseteq$ inclusion, let $f\in\mathcal{H}_{a}$. Then there is an open finite-index subgroup $V\leq G_{a}$ which fixes $f$. If we suppose that there is $b\in M\setminus\operatorname{acl}a$ such that $f(b)\neq 0$, we have that $f$ is constant on $V\cdot b$, which is infinite, and this contradicts the fact that $f\in\ell^{2}$. Conversely, write $\operatorname{acl}a=\bigcup_{i}A_{i}$, where each $A_{i}$ is finite, $G_{a}$-invariant, and $A_{i}\subseteq A_{i+1}$. Then for each $f$ with $\operatorname{supp}f\subseteq\operatorname{acl}a$, we have that $f\chi_{A_{i}}\to f$ and $G_{a}\cdot(f\chi_{A_{i}})$ is finite for each $i$. Now, recalling that $\operatorname{acl}(a\wedge b)=\operatorname{acl}a\cap\operatorname{acl}b$, the conclusion is clear for $\mathcal{H}=\ell^{2}(\mathcal{U})$.
@@ 3.1. Non-singular actions
An action on $\mathrm{MALG}(\mu)$ as above is often called in the literature a Boolean action. It is clear that any pointwise Borel action of $G$ on $\Omega$ such that $g_{*}\mu\sim\mu$ for all $g\in G$ gives rise to a Boolean action. The converse, for pmp actions and non-archimedean Polish groups, is also true, as shown by Glasner and Weiss in [Glasner2005a]*Theorem 2.3. Subsection 3.3 below will imply that the same holds for non-singular actions of Roelcke precompact, non-archimedean, Polish groups.
@@
    Suppose that $f_{0}\in L^{2}(\Omega)$ with $\lVert f_{0}\rVert=1$ is fixed. Then $|f_{0}|$ is also fixed and positive of norm $1$. Let $\nu$ be the probability measure defined by $\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}=|f_{0}|^{2}$. Then, for every $g\in G$, using the equation (missing) 3.1, we have
@@
    One checks using the same calculation as above that the element $\big{(}\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}\big{)}^{1/2}\in L^{2}(\Omega)$ is invariant. ∎

### v2

@@ 1. Introduction
Roelcke precompact groups share some tameness properties with compact groups. For example, the unitary representations of non-archimedean Roelcke precompact groups can be completely classified [Tsankov2012]. However, in contrast to the situation with compact groups, irreducible representations are usually infinite-dimensional. This classification has proved helpful for studying their probability measure-preserving (pmp, for short) actions. For example, an analogue of de Finetti’s theorem holds for appropriate actions of such groups ([Jahel2022], also see below) and Jahel and Joseph [Jahel2025] have been able to classify the invariant random subgroups of some of them.
@@
Let $G$ be a non-archimedean, Roelcke precompact, Polish group and let $G\curvearrowright(\Omega,\mu)$ be a non-singular action. Then $\Omega$ can be represented as a disjoint countable union $\bigsqcup_{i}\Omega_{i}$ of invariant subsets on each of which the action is isomorphic to an induced action from a pmp action of an open subgroup of $G$. In particular, there exists a $\sigma$-finite measure equivalent to $\mu$ which is preserved by $G$.
@@
Then any $G$-invariant, ergodic measure on $0,1]^{M}$ is a product measure.
Condition (ii) above is satisfied, for example, if the action $G\curvearrowright M^{2}$ has finitely many orbits but also in many other situations. This condition has been considered in the pseudo-finite case in [Liebeck2010]. Both conditions (i) and (ii) are remnants of compactness and are automatically satisfied if $G$ is Roelcke precompact. Condition (i) is also satisfied, for example, if $M$ is a homogeneous structure whose age is given by a first-order universal theory (or, equivalently, a collection of finite forbidden configurations) and $G=\operatorname{Aut}(M)$. An important class of examples where the hypotheses are not satisfied but the conclusion still holds is the rational Urysohn space and some generalizations of it. In a very recent preprint [Barritault2024pp], Barritault, Jahel, and Joseph manage to classify all unitary representations of its isometry group and deduce, in particular, a version of Section 1.
@@ Acknowledgments
I would like to thank Tomás Ibarlucía and Yuri Neretin for bringing to my attention Nessonov’s article [Nessonov2020], which inspired Section 1. I am also grateful to Rémi Barritault, Colin Jahel, and Matthieu Joseph for useful remarks on a preliminary version of the paper and to the anonymous referee for useful suggestions.
@@ 2.1. The lattice of algebraically closed sets
(i) Write $H=G_{D}$. Let $S_{0},S_{1},\ldots$ be an enumeration of the orbits of $G$ on $M$. Let $A_{i}=\bigcup_{j\leq i}S_{j}\cap D$. As $D$ is locally finite, each $A_{i}$ is finite. Moreover, $G_{A_{0}}\supseteq G_{A_{1}}\supseteq\cdots$ and $\bigcap_{i}G_{A_{i}}=H$. As $H$ is open, by Roelcke precompactness, each $G_{A_{i}}$ contains only finitely many $H$-double cosets, so there is $i$ such that $G_{A_{i}}=H$. Let $A=A_{i}$. We check that $\operatorname{acl}A=D$. The $\subseteq$ inclusion follows from the fact that $D$ is algebraically closed. For the other, if $a\in D$, then $G_{A}\cdot a=H\cdot a\subseteq D\cap G\cdot a$, which is finite by the local finiteness of $D$.
(ii) If $gG_{A}g^{-1}\cap G_{A}$ has finite index in $G_{A}$, then $G_{A}g\cdot A$ is finite. Thus $g\cdot A\subseteq\operatorname{acl}A=D$, which implies that $g\cdot A\subseteq A$. As $A$ is finite, this means that $g\cdot A=A$ and thus $g\in G_{A}=G_{D}$.
(iii) This follows directly from (ii). ∎
@@
(i) This follows from [Evans2016]*Lemma 2.4, noting that for $K\leq G$ and $A\in\operatorname{Fin}(M)$, the orbit $K\cdot A$ is infinite iff there exists $a\in A$ with $K\cdot a$ infinite.
(ii) Note that $H\lhd G_{D}$ and that $G_{D}/H$ acts faithfully on $D$ with finite orbits. It follows from [Evans2016]*Lemma 2.9 that the action $G_{D}/H\curvearrowright D$ is topologically faithful, so $G_{D}/H$ is a profinite group. As every $G$-orbit splits into finitely many $G_{D}$-orbits by Roelcke precompactness, it suffices to show that every $G_{D}$-orbit splits into finitely many $H$-orbits. Let $A\in\operatorname{Fin}(M)$ and $F\subseteq G_{D}$ be finite such that $G_{D}=HF(G_{A}\cap G_{D})$. Then $F\cdot A$ is a set of representatives for all $H$-orbits in $G_{D}\cdot A$. ∎
@@
At the level of stabilizers, $A\leq B$ iff $G_{A}\cap G_{B}$ has finite index in $G_{B}$, and $A\sim B$ iff $G_{A}$ and $G_{B}$ are commensurate. It follows from () 2.1 that for all $A,B\in\operatorname{Fin}(M)$,
@@
Let $G$ be a non-archimedean, Roelcke precompact, Polish group and let $G\curvearrowright M$ be a continuous action on a discrete, countable set. Let the quasi-order $\leq$ be defined as in () 2.1. Then the following hold:
@@
    The relation << on $\operatorname{Fin}(M)$ is well-founded.
@@
(i) Let $A,B\in\operatorname{Fin}(M)$. The least upper bound $A]\vee[B]$ is given by $A\cup B]$. That $A\cup B]$ is an upper bound follows from () 2.2. If $C]\geq[A],[B]$, then $A\cup B\subseteq\operatorname{acl}C$, so $\operatorname{acl}(A\cup B)\subseteq\operatorname{acl}C$ by Subsection 2.1.
For the greatest lower bound, let $D=\operatorname{acl}A\cap\operatorname{acl}B$ and note that $D$ is algebraically closed and locally finite. It is also definable because $G_{(A\cup B)}$ is a subgroup of $G_{D}$. Use Subsection 2.1 to find a finite $D_{0}\subseteq D$ with $D=\operatorname{acl}D_{0}$ and $G_{D}=G_{D_{0}}$. We show that $D_{0}]=[A]\wedge[B]$. It is clear that $D_{0}]\leq[A],[B]$. If $C\in\operatorname{Fin}(M)$ is such that $C\leq A,B$, then $C\subseteq\operatorname{acl}A\cap\operatorname{acl}B$, so $\operatorname{acl}C\subseteq\operatorname{acl}A\cap\operatorname{acl}B=\operatorname{acl}D_{0}$, and it follows that $C\leq D_{0}$.
(ii) Suppose, towards a contradiction, that there exist $A_{0},A_{1},\ldots\in\operatorname{Fin}(M)$ with $A_{0}>A_{1}>\cdots$. Let $D_{i}=\operatorname{acl}A_{i}$ and note that $D_{i}\supseteq D_{i+1}$ and $A_{i}\subseteq D_{j}$ for all $i\geq j$. Let $H_{i}=G_{(D_{i})}$, so that $H_{0}\leq H_{1}\leq\cdots$.
We prove that for every $i$, the inclusion $H_{i}\cdot A_{0}\subseteq H_{i+1}\cdot A_{0}$ is proper. By assumption, $G_{A_{i+1}}\cdot A_{i}$ is infinite, so $H_{i+1}\cdot A_{i}$ is also infinite by Subsection 2.1 (i). On the other hand, $A_{i}\subseteq\operatorname{acl}A_{0}$, so $G_{A_{0}}\cdot A_{i}$ is finite. Hence there exists $g\in H_{i+1}$ with $g^{-1}\cdot A_{i}\notin G_{A_{0}}\cdot A_{i}$. We will show that $g\cdot A_{0}\notin H_{i}\cdot A_{0}$. Suppose, to the contrary, that there is $h\in H_{i}$ with $h\cdot A_{0}=g\cdot A_{0}$. Then $g^{-1}h\in G_{A_{0}}$ and $g^{-1}h\cdot A_{i}=g^{-1}\cdot A_{i}$, contradicting the fact that $g^{-1}\cdot A_{i}\notin G_{A_{0}}\cdot A_{i}$.
We obtained that $H_{0}\cdot A_{0}\subsetneq H_{1}\cdot A_{0}\subsetneq\cdots$ is an infinite strictly increasing sequence of $H_{0}$-invariant subsets of $G\cdot A_{0}$, which contradicts Subsection 2.1 (ii). ∎
@@ 2.2. The universal action and unitary representations
Let $G$ be a Roelcke precompact, non-archimedean, Polish group. Among all actions of $G$ on a countable set, there is a universal one that can be constructed as follows. Let $(V_{i}:i\in I)$ be a collection of representatives for the equivalence relation of conjugacy on open subgroups of $G$. Note that the set $I$ is countable by [Tsankov2012]*Corollary 2.5. Let
@@
and equip it with the left translation action. Note that for any open $V\leq G$, there exists $a\in\mathcal{U}$ with $V=G_{a}$. This implies, in particular, that the action $G\curvearrowright\mathcal{U}$ is topologically faithful (see [Evans2016]*Lemma 1.9). This action is universal in the sense that every continuous, transitive action of $G$ on a countable set embeds into it and it is minimal with this property. If one wants to embed any continuous action of $G$ on a countable set, one has to take infinitely many copies of $\mathcal{U}$.
@@
Indeed, by Subsection 2.1 and Subsection 2.1, $G_{\operatorname{acl}A}$ is an open subgroup of $G$. Let $V_{i}$ be the representative of its conjugacy class that appears in the definition of $\mathcal{U}$ and let $g\in G$ be such that $gV_{i}g^{-1}=G_{\operatorname{acl}A}$. Then one can take $a=gV_{i}$. To verify uniqueness, note that $G_{gV_{i}}=gV_{i}g^{-1}$. By the construction of $\mathcal{U}$, if $G_{gV_{i}}=G_{hV_{j}}$, then $V_{i}=V_{j}$ and $h^{-1}g\in N_{G}(V_{i})$. By Subsection 2.1, $V_{i}$ is self-normalizing, so $hV_{i}=gV_{i}$.
It is easy to check that $A\sim s(A)$ for all $A\in\operatorname{Fin}(\mathcal{U})$ and that $s$ is idempotent in the sense that $s(\{s(A)\})=s(A)$ for all $A\in\operatorname{Fin}(M)$. Thus the set
@@
Notice that the subspace on the right-hand side is closed. To see the $\subseteq$ inclusion, let $f\in\mathcal{H}_{a}$ be such that $G_{a}\cdot f$ is finite. Then there is an open finite-index subgroup $V\leq G_{a}$ which fixes $f$. If we suppose that there is $b\in M\setminus\operatorname{acl}a$ such that $f(b)\neq 0$, we have that $f$ is constant on $V\cdot b$, which is infinite, and this contradicts the fact that $f\in\ell^{2}(\mathcal{U})$. Conversely, write $\operatorname{acl}a=\bigcup_{i}A_{i}$, where each $A_{i}$ is finite, $G_{a}$-invariant, and $A_{i}\subseteq A_{i+1}$. Then for each $f$ with $\operatorname{supp}f\subseteq\operatorname{acl}a$, we have that $f\chi_{A_{i}}\to f$ and $G_{a}\cdot(f\chi_{A_{i}})$ is finite for each $i$. Now, recalling that $\operatorname{acl}(a\wedge b)=\operatorname{acl}a\cap\operatorname{acl}b$, the conclusion is clear for $\mathcal{H}=\ell^{2}(\mathcal{U})$.
@@ 3.1. Non-singular actions
An action on $\mathrm{MALG}(\mu)$ as above is often called in the literature a Boolean action. It is clear that any pointwise Borel action of $G$ on $\Omega$ such that $g_{*}\mu\sim\mu$ for all $g\in G$ gives rise to a Boolean action. The converse, for pmp actions and non-archimedean Polish groups, is also true, as shown by Glasner and Weiss in [Glasner2005a]*Theorem 2.3. Subsection 3.3 below implies that the same holds for non-singular actions of Roelcke precompact, non-archimedean, Polish groups.
@@
    Suppose that $f_{0}\in L^{2}(\Omega)$ with $\lVert f_{0}\rVert=1$ is fixed. Then $|f_{0}|$ is also fixed and positive of norm $1$. Let $\nu$ be the probability measure defined by $\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}=|f_{0}|^{2}$. Then, for every $g\in G$, using the equation () 3.1, we have
@@
    We will check that $\big(\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}\big)^{1/2}\in L^{2}(\Omega)$ is invariant. Indeed,
    $\begin{split}\Big(g\cdot\Big(\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}\Big)^{1/2}\Big)(x)&=\Big(\frac{\,\mathrm{d}g_{*}\mu}{\,\mathrm{d}\mu}\Big)^{1/2}(x)\Big(\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}\Big)^{1/2}(g^{-1}\cdot x)\\ &=\Big(\frac{\,\mathrm{d}g_{*}\mu}{\,\mathrm{d}\mu}\Big)^{1/2}(x)\Big(\frac{\,\mathrm{d}g_{*}\nu}{\,\mathrm{d}g_{*}\mu}\Big)^{1/2}(x)\\ &=\Big(\frac{\,\mathrm{d}g_{*}\mu}{\,\mathrm{d}\mu}\Big)^{1/2}(x)\Big(\frac{\,\mathrm{d}\nu}{\,\mathrm{d}g_{*}\mu}\Big)^{1/2}(x)=\Big(\frac{\,\mathrm{d}\nu}{\,\mathrm{d}\mu}\Big)^{1/2}(x).\qed\end{split}$